OpenStudy (anonymous):

h(x)=x^2/x^3-4x...how do you find the domain

OpenStudy (anonymous):

the domain is all x's that the equation is defined for example the domain of x^2 is all real numbers because you can put anything in for x the domain of 1/x is all reals except zero because you cannot divide by zero what values can't you plug into the original equation? (things that would result in divide by zero)

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

example the function 1/(x+3) what number could you plug in to make the denominator equal to zero?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

right so -3 is not in the domain all other numbers are

OpenStudy (anonymous):

i have (-00,-2)U(2,00)..thats da infinity symbol

OpenStudy (anonymous):

so (-3,00)?

OpenStudy (anonymous):

and everything on the other side too (-00,-3)U(-3,00)

OpenStudy (anonymous):

ok so mines isnt rite for my problem

OpenStudy (anonymous):

is the problem like this (x^2/x^3)-4x or like this (x^2)/(x^3-4x)

OpenStudy (anonymous):

(x-2)/(x^3-4x)

OpenStudy (anonymous):

ok so what numbers would make the denominator zero?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

and -2

OpenStudy (anonymous):

and 0

OpenStudy (anonymous):

so the domain is everything except those 3 numbers

OpenStudy (anonymous):

(-00,-2)U(-2,0,2)U(2,00)?

OpenStudy (anonymous):

close check your notation the middle part should be (-2,0)U(0,2)

OpenStudy (anonymous):

why would you write it that way

OpenStudy (anonymous):

and one more question

OpenStudy (anonymous):

to answer your first question, its because 0 is not included, just like -2 and 2 you have to write it that way to show that 0 isnt included

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what's your other question?

OpenStudy (anonymous):

how do you graph g(x)=(-√x+2)+2 without usin a calculator

OpenStudy (anonymous):

is the first 2 under the square root?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

in that case, its really easy do you know what the graph of y=sqrt(x) looks like?

OpenStudy (anonymous):

yea but i dont know how you tie that in to the problem like how do you know which way the graph moves and wheres its suppose to be

OpenStudy (anonymous):

y=sqrt(x) substitute in x+2 for x moves the entire graph left 2 spaces ****IMPORTANT***** this wouldnt be the case if it was y=sqrt(x)+2 it has to be y=sqrt(x+2) if the number with x is positive, the graph is moved to the left if its negative, its moved to the right the number outside sqrt(x) tells how far up or down to move it example sqrt(x)+1 is moved up 1 if it was -1 instead, it would be moved down 1 a negative in front of the entire thing mirrors it over the x axis

OpenStudy (anonymous):

back to the original since the main part of the function is negative, it gets flipped over the x axis a substitution of x+2 was made for x (its all in the parenthesis) so the graph gets moved left 2 spaces 2 was added outside the parenthesis/main part) so the graph is moved up 2 spaces ill post a drawing

OpenStudy (anonymous):